# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

### v1

@@ Abstract
Spiking Neural Networks (SNNs), as a biologically plausible alternative to Artificial Neural Networks (ANNs), have demonstrated advantages in terms of energy efficiency, temporal processing, and biological plausibility. However, SNNs are highly sensitive to distribution shifts, which can significantly degrade their performance in real-world scenarios. Traditional test-time adaptation (TTA) methods designed for ANNs often fail to address the unique computational dynamics of SNNs, such as sparsity and temporal spiking behavior. To address these challenges, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method specifically designed for SNNs. SPACE leverages the inherent spike dynamics of SNNs to maximize the consistency of spike-behavior-based local feature maps across augmented versions of a single test sample, enabling robust adaptation without requiring source data. We evaluate SPACE on multiple datasets, including CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C and DVS Gesture-C. Furthermore, SPACE demonstrates strong generalization across different model architectures, achieving consistent performance improvements on both VGG9 and ResNet11. Experimental results show that SPACE outperforms state-of-the-art methods, highlighting its effectiveness and robustness in real-world settings.
@@ 1 Introduction
    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [24], Tiny-ImageNet (a modified subset of the ImageNet dataset) [6], and the neuromorphic dataset DVS Gesture [1]. Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on VGG9 [37] and ResNet11 [14]. The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
@@ 2.1 SNN-based Deep Learning
Over the years, significant progress has been made in SNN-based deep learning. Many approaches primarily relied on converting pre-trained ANN models into SNNs, enabling SNNs to inherit the representational power of ANNs while benefiting from spike-based computations [3, 8, 17, 20]. However, such conversion methods often suffer from performance degradation due to differences in activation dynamics. To overcome this, direct training of SNNs using surrogate gradient methods has gained traction, enabling end-to-end optimization of spiking models [7, 11, 29, 31, 42]. Rathi et al. [33] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance. These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [26, 40] and spiking recurrent networks [41, 44], which have demonstrated promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
A critical limitation of current SNN-based deep learning approaches is their poor robustness to testing-time variations. Unlike ANNs, which operate on continuous activations and can leverage well-established testing-time adaptation techniques, SNNs are highly sensitive to input perturbations due to their dependence on precise spike timing and temporal dynamics. According to our experimental results presented in Sec. 4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time. Addressing these shortcomings is essential for deploying SNNs in practical applications, particularly those requiring real-time adaptability.
@@ 3.1 Preliminary
where $U(t)$ represents the membrane potential of the neuron at time $t$, $\tau_{m}$ is the membrane time constant, $R$ denotes the input resistance and $I(t)$ is the input current received from pre-synaptic neurons or inputs. When $U(t)$ exceeds a predefined threshold $U_{th}$, the neuron emits a spike and its membrane potential is reset to a resting value, typically 0 or $U(t)-U_{th}$. Following [11], Eq. 1 is discretized as
@@
Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for weight connections between pre- and post-neurons. Fig. 1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
@@ 3.2 SPike-Aware Consistency Enhancement
Figure 2: Overview of the SPike-Aware Consistency Enhancement (SPACE) framework for TTA in SNNs. The test sample is selected from CIFAR-10-C dataset [15] with Gaussian Noise corruption at level 5. The model here follows the VGG9 [37] architecture. The process involves four main steps: 1) Generate an augmented batch from the single test sample. 2) Pass the augmented batch through the model to obtain local feature maps, represented by the spike counts over the processing time. 3) Adapt the model by maximizing the similarity across the local feature maps of the augmented samples. 4) Use the adapted model to predict the label of the original test sample.
@@
Algorithm Design Overview. The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Fig. 2. We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample. This augmented batch is passed through the model to obtain local feature maps, represented by spike counts over time. Next, the model is adapted by maximizing the similarity across the feature maps of the augmented samples, promoting consistency in feature representations. Finally, the adapted model predicts the label of the original test sample, ensuring robust performance under test-time conditions. This approach enhances the model’s ability to generalize to unseen test samples, particularly in shifted domains. The overall method is presented in Algorithm 1, and the details are introduced below. The code will be made available after acceptance.
@@
1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, SNN model with parameters $\theta$, and learning rate $\eta$
2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the adapted model
@@
4:  Initialize augmented batch $\mathcal{B}=\emptyset$
5:  for each $a_{k}$ in subset $\{a_{1},\dots,a_{M}\}$ of $\mathcal{A}$ do
6:     Apply augmentation: $\mathbf{x}_{k}=a_{k}(\mathbf{x})$
7:     Add $\mathbf{x}_{k}$ to $\mathcal{B}$
8:  end for
9:  Step 2: Extract Feature Maps
10:  Pass $\mathcal{B}$ through the SNN feature extractor
11:  Obtain feature maps:
12:  =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
13:  Step 3: Compute Similarity
14:  for each pair $(\mathbf{F}(\mathbf{x}_{i}),\mathbf{F}(\mathbf{x}_{j}))$ in $\mathcal{F}$ do
15:     Compute similarity $\bar{\mathcal{S}}(i,j|\mathbf{x})$ using Eq. 4 $\sim$ Eq. 6
16:  end for
17:  Step 4: Compute loss and update model
@@
18:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Eq. 7
19:  Update parameters using Eq. 8
20:  Step 5: Predict via adapted model
21:  Predict via: $y^{*}=argmax_{y}p_{\theta^{*}}(y|\mathbf{x})$
22:  Return: Prediction result $y^{*}$
Spike-Aware Feature Maps. SPACE aims to adapt pre-trained SNN-based models $M_{\theta}$ (with parameters $\theta\in\Theta$) during inference to improve performance under distribution shifts, without requiring ground truth labels or access to source data. No specific training process is required, nor do we impose particular constraints on the model. The only assumptions are that $\theta$ can be adjusted and that the model produces intermediate feature maps $\mathbf{F}(\mathbf{x})$, which is differentiable with respect to $\theta$ and can be utilized for further analysis and adaptation. Here, $\mathbf{x}\in\mathcal{X}$ is a single test point that is presented to $M_{\theta}$. It is worth noting that this is a reasonable assumption, supported by a recent advance [34] that suggest gradient back-propagation can be implemented on spiking neuromorphic hardware. Aligned with previous work [21, 43], we achieve test-time robustness by leveraging data augmentations and a self-supervised adaptation objective. In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples =$\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$. Using this batch, we align the spike dynamics extracted from augmentations, ensuring that the model produces reliable and robust predictions in the presence of distribution shifts.
To leverage the feature maps from the intermediate layers of the SNN, we focus on the spike activity over the entire time window. As described in Sec. 3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$. The temporal spike patterns of the network are captured in the feature map $\mathbf{O(\mathbf{x}_{i})}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the number of time steps, $C$ represents the number of channels, and $H$ and $W$ are the spatial dimensions of the local feature maps. $\mathbf{x}_{i}$ refers to the $i$-th sample in the augmented batch. Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network (typically a CNN [25]), depending on whether the final feature map is pooled into a 1$\times$1 matrix. Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
@@
Feature Maps Alignment. The primary objective is to encourage the model to focus on invariant features that are crucial for classification, rather than being influenced by noise or minor transformations introduced by augmentations or distribution shifts in the training data. When the feature maps of augmented samples are highly similar, the model becomes less sensitive to variations at test time, thereby improving its ability to generalize to new, unseen, and unlabeled data. To measure the similarity between feature representations of the $i$-th and $j$-th samples, we employ channel-wise similarity between corresponding local feature vectors, $\mathbf{F}_{c}(\mathbf{x}_{i})$ and $\mathbf{F}_{c}(\mathbf{x}_{j})$. Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution:
$\mathbf{P}_{c}(\mathbf{x}_{i})=\textit{softmax}(\mathbf{F}_{c}(\mathbf{x}_{i})).$ (4)
Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ represent the local feature vectors corresponding to channel $c$ of the augmented sample $\mathbf{x}_{i}$. This normalization highlights important spatial locations by amplifying prominent feature values and suppressing noise, enhancing feature distribution stability while enabling the subsequent similarity computation to capture both spatial and temporal characteristics effectively. To quantify the similarity between $\mathbf{F}(\mathbf{x}_{i})$ and $\mathbf{F}(\mathbf{x}_{j})$, we compute the average of channel-wise weighted inner product, defined as
$\mathcal{S}_{c}(i,j|\mathbf{x})=\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (5)
$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\mathcal{S}_{c}(i,j|\mathbf{x}).$ (6)
@@
Then we adapt the model via gradient descent for only one iteration given as
$\theta^{*}\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta;\mathbf{x}),$ (8)
where $\eta$ is learning rate here. By minimizing this loss function, we enforce consistency across the feature maps of different augmentations. This reduces the variability induced by augmentations, ensuring that the model learns stable and robust feature representations. Furthermore, by promoting feature consistency, this approach serves as a regularization mechanism, preventing over-fitting to specific augmentations and encouraging the model to focus on more fundamental and generalizable features.
While the proposed similarity measure effectively captures feature distribution overlap, we further investigated whether incorporating higher-dimensional feature relationships could enhance the consistency measure. Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) (see Sec. 6 in Supplementary Material). However, this method did not yield noticeable performance improvements and introduced additional computational overhead, suggesting that the original similarity measure is already sufficient for our test-time adaptation.
@@
Discussion. In SNNs, each local feature map captures distinct aspects of the input data and exhibits different spike patterns, both spatially and temporally. Since channels may have different active neurons, directly comparing the entire feature map fails to account for this variability. Instead, computing the similarity per channel and averaging the results focuses on the stability of feature representations within each channel. This isolates the contributions of active neurons, ensuring the similarity measure reflects meaningful changes rather than noise. It also minimizes the impact of irregular spike behavior or augmentation-induced noise in individual channels. Our experiments in Sec. 4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach. Given that each channel learns distinct features, averaging the similarity within each channel provides a more reliable measure of feature map consistency, improving TTA performance.

### v2

@@ Abstract
Spiking Neural Networks (SNNs), as a biologically plausible alternative to Artificial Neural Networks (ANNs), have demonstrated advantages in terms of energy efficiency, temporal processing, and biological plausibility. However, SNNs are highly sensitive to distribution shifts, which can significantly degrade their performance in real-world scenarios. Traditional test-time adaptation (TTA) methods designed for ANNs often fail to address the unique computational dynamics of SNNs, such as sparsity and temporal spiking behavior. To address these challenges, we propose SPike-Aware Consistency Enhancement (SPACE), the first source-free and single-instance TTA method specifically designed for SNNs. SPACE leverages the inherent spike dynamics of SNNs to maximize the consistency of spike-behavior-based local feature maps across augmented versions of a single test sample, enabling robust adaptation without requiring source data. We evaluate SPACE on multiple datasets. Furthermore, SPACE exhibits robust generalization across diverse network architectures, consistently enhancing the performance of SNNs on CNNs (such as VGG and ResNet), Transformer models, and ConvLSTM architectures. Experimental results show that SPACE outperforms state-of-the-art methods, highlighting its effectiveness and robustness in real-world settings.
@@ 1 Introduction
    To assess the effectiveness and robustness of our proposed method, we perform extensive experiments across multiple benchmarks, including CIFAR-10, CIFAR-100 [26], Tiny-ImageNet, ImageNet [6], and the neuromorphic dataset DVS Gesture [1]. Furthermore, we validate the adaptability of our method across different model architectures, specifically testing on SNN-VGG [24, 40], SNN-ResNet11 [14, 27], Spike-driven Transformer V3 [44], and SNN-ConvLSTM [39]. The results demonstrate that our approach not only achieves consistent performance improvements across datasets but also generalizes well to different network structures, showcasing its broad applicability and robustness.
@@ 2.1 SNN-based Deep Learning
Over the years, significant progress has been made in SNN-based deep learning. Many approaches primarily relied on converting pre-trained ANN models into SNNs, enabling SNNs to inherit the representational power of ANNs while benefiting from spike-based computations [3, 8, 19, 22]. However, such conversion methods often suffer from performance degradation due to differences in activation dynamics. To overcome this, direct training of SNNs using surrogate gradient methods has gained traction, enabling end-to-end optimization of spiking models [7, 11, 30, 32, 46]. Rathi et al. [34] proposed a hybrid method that combines conversion and surrogate gradient-based method, achieving state-of-the-art performance. These advancements have not only solved the challenge owing to the non-differentiable nature of spike functions and the the inherent complexity of temporal dynamics, but also led to the development of sophisticated SNN architectures, such as spiking convolutional networks [27, 43] and spiking recurrent networks [45, 48], which have denoted promising results in tasks like image classification, speech recognition, and event-driven sensor data analysis.
A critical limitation of current SNN-based deep learning approaches is their poor robustness to testing-time variations. Unlike ANNs, which operate on continuous activations and can leverage well-established testing-time adaptation techniques, SNNs are highly sensitive to input perturbations due to their dependence on precise spike timing and temporal dynamics. According to our experimental results presented in section 4, existing SNN models, which are typically trained offline, lack mechanisms to dynamically adapt to changing environments or incoming data distributions at testing time. Addressing these shortcomings is essential for deploying SNNs in practical applications, particularly those requiring real-time adaptability.
@@ 3.1 Preliminary
where $U(t)$ represents the membrane potential of the neuron at time $t$, $\tau_{m}$ is the membrane time constant, $R$ denotes the input resistance and $I(t)$ is the input current received from pre-synaptic neurons or inputs. When $U(t)$ exceeds a predefined threshold $U_{th}$, the neuron emits a spike and its membrane potential is reset to a resting value, typically 0 or $U(t)-U_{th}$. Following [11], Equation 1 is discretized as
@@
Here, $j$ is the index of pre-synaptic neurons, $o_{j}$ is the binary spike activation, and $w_{ij}$ stands for weight connections between pre- and post-neurons. Figure 1 illustrates the LIF neuron model, depicting membrane potential evolution and spike generation in response to input spikes [10].
@@ 3.2 SPike-Aware Consistency Enhancement
Figure 2: Overview of the SPike-Aware Consistency Enhancement (SPACE) framework for TTA in SNNs. The test sample is selected from CIFAR-10-C dataset [15] with Gaussian Noise corruption at level 5. The model here follows the VGG9 architecture. The process involves four main steps: 1) Generate an augmented batch from the single test sample. 2) Pass the augmented batch through the model to obtain local feature maps, represented by the spike counts over the processing time. 3) Adapt the model by maximizing the similarity across the local feature maps of the augmented samples. 4) Use the adapted model to predict the label of the original test sample.
Algorithm Design Overview
@@
The proposed algorithm through spike-aware consistency enhancement (SPACE) for test-time adaptation is designed to improve test-time robustness of SNNs, as shown in Figure 2. We first generate an augmented batch from a single test sample using various augmentation techniques, introducing diversity while retaining the core characteristics of the original sample. This augmented batch is passed through the model to obtain local feature maps, represented by spike counts over time. Next, the model is adapted by maximizing the similarity across the feature maps of the augmented samples, promoting consistency in feature representations. Finally, the adapted model predicts the label of the original test sample, ensuring robust performance under test-time conditions. This approach enhances the model’s ability to generalize to unseen test samples, particularly in shifted domains. The overall method is presented in Algorithm 1, and the details are introduced below. The code will be made available after acceptance.
Spike-Aware Feature Maps
SPACE aims to adapt pre-trained SNN-based models $M_{\theta}$ (with parameters $\theta\in\Theta$) during inference to improve performance under distribution shifts, without requiring ground truth labels or access to source data. No specific training process is required, nor do we impose particular constraints on the model. The only assumptions are that $\theta$ can be adjusted and that the model produces intermediate feature maps $\mathbf{F}(\mathbf{x})$, which is differentiable with respect to $\theta$ and can be utilized for further analysis and adaptation. Here, $\mathbf{x}\in\mathcal{X}$ is a single test point that is presented to $M_{\theta}$. It is worth noting that this is a reasonable assumption, supported by a recent advance [36] that suggest gradient back-propagation can be implemented on spiking neuromorphic hardware. Aligned with previous work [23, 47], we achieve test-time robustness by leveraging data augmentations and a self-supervised adaptation objective. In our work, a set of augmentation functions with different intensities selected from $\mathcal{A}\triangleq\{a_{1},\dots,a_{N}\}$ is applied to a single test input $\mathbf{x}$, forming an augmented batch of samples =$\mathcal{B}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{M}\},M\leq N$. Using this batch, we align the spike dynamics extracted from augmentations, ensuring that the model produces reliable and robust predictions in the presence of distribution shifts.
To leverage the feature maps from the intermediate layers of the SNN, we focus on the spike activity over the entire time window. As described in subsection 3.1, each neuron $j$ in an SNN can either fire a spike (outputting 1) or remain silent (outputting 0) at each time step $t$, represented by the binary function $o_{j}^{t}$. The temporal spike patterns of the network are captured in the feature map $\mathbf{O}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the number of time steps, $C$ represents the number of channels, and $H$ and $W$ are the spatial dimensions of the local feature maps. $\mathbf{x}_{i}$ refers to the $i$-th sample in the augmented batch. Our primary objective is to use the total spike counts of the feature maps from either the last layer or the penultimate layer of the feature extraction network, depending on whether the final feature map is pooled into a 1$\times$1 matrix. Let =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\mathbf{F}(\mathbf{x}_{2}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$ denote the collection of feature maps, where each $\mathbf{F}(\mathbf{x}_{i})\in\mathbb{R}^{C\times D}$ represents the total spike counts of neurons across different channels, with $D=H\times W$ indicating the spatial dimensionality.
@@
1:  Input: Test sample $\mathbf{x}$, augmentation functions $\mathcal{A}=\{a_{1},\dots,a_{N}\}$, model $\theta$, learning rate $\eta$
2:  Output: Prediction $y^{*}$ of $\mathbf{x}$ using the updated model
@@
4:  Initialize $\mathcal{B}=\emptyset$
5:  for $a_{k}$ in a subset $\{a_{1},\dots,a_{M}\}$ of $\mathcal{A}$ do
6:     $\mathbf{x}_{k}=a_{k}(\mathbf{x})$, add $\mathbf{x}_{k}$ to $\mathcal{B}$
7:  end for
8:  Step 2: Extract feature maps
9:  Pass $\mathcal{B}$ through SNN, obtain =$\mathcal{F}=\{\mathbf{F}(\mathbf{x}_{1}),\dots,\mathbf{F}(\mathbf{x}_{M})\}$
10:  Step 3: Compute similarity
11:  for each pair of ($\mathbf{F}(\mathbf{x}_{i}),\mathbf{F}(\mathbf{x}_{j})$) in $\mathcal{F}$ do
12:     Compute similarity $\bar{\mathcal{S}}(i,j|\mathbf{x})$ using Equation 4
13:  end for
14:  Step 4: Update model parameters
@@
15:  Compute loss $\mathcal{L}(\theta;\mathbf{x})$ using Equation 5, update $\theta$ via SGD with $\eta$
16:  Step 5: Predict
17:  $y^{*}=\arg\max_{y}p_{\theta^{*}}(y|\mathbf{x})$
18:  Return: $y^{*}$
Feature Maps Alignment
@@
The primary objective is to encourage the model to focus on invariant features that are crucial for classification, rather than being influenced by noise or minor transformations introduced by augmentations or distribution shifts in the training data. When the feature maps of augmented samples are highly similar, the model becomes less sensitive to variations at test time, thereby improving its ability to generalize to new, unseen, and unlabeled data. To measure the similarity between feature representations of the $i$-th and $j$-th samples, we employ channel-wise similarity between corresponding local feature vectors, $\mathbf{F}_{c}(\mathbf{x}_{i})$ and $\mathbf{F}_{c}(\mathbf{x}_{j})$. Here, $\mathbf{F}_{c}(\mathbf{x}_{i})$ denotes the local feature vectors associated with channel $c$ of the augmented sample $\mathbf{x}_{i}$. Specifically, we first normalize local feature vectors using a softmax function along the spatial dimensions, ensuring that the features are transformed into a probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$. This normalization highlights important spatial locations by amplifying prominent feature values and suppressing noise, enhancing feature distribution stability while enabling the subsequent similarity computation to capture both spatial and temporal characteristics effectively. To quantify the similarity between $\mathbf{F}(\mathbf{x}_{i})$ and $\mathbf{F}(\mathbf{x}_{j})$, we compute the average of channel-wise weighted inner product, defined as
$\bar{\mathcal{S}}(i,j|\mathbf{x})=\frac{1}{C}\sum_{c=1}^{C}\sum_{d=1}^{D}\mathbf{P}_{c,d}(\mathbf{x}_{i})\cdot\mathbf{P}_{c,d}(\mathbf{x}_{j}),$ (4)
@@
Then, we adapt the model using stochastic gradient descent (SGD) for a single iteration with the learning rate $\eta$. By minimizing this loss function, we enforce consistency across the feature maps of different augmentations. This reduces the variability induced by augmentations, ensuring that the model learns stable and robust feature representations. Furthermore, by promoting feature consistency, this approach serves as a regularization mechanism, preventing over-fitting to specific augmentations and encouraging the model to focus on more fundamental and generalizable features.
While the proposed similarity measure effectively captures feature distribution overlap, we further investigated whether incorporating higher-dimensional feature relationships could enhance the consistency measure. Specifically, we experimented with a kernel embedding approach combined with Maximum Mean Discrepancy (MMD) in Appendix). However, this method did not yield noticeable performance improvements and introduced additional computational overhead, suggesting that the original similarity measure is already sufficient for our test-time adaptation.
@@
Discussion
In SNNs, each local feature map captures distinct aspects of the input data and exhibits different spike patterns, both spatially and temporally. Since channels may have different active neurons, directly comparing the entire feature map fails to account for this variability. Instead, computing the similarity per channel and averaging the results focuses on the stability of feature representations within each channel. This isolates the contributions of active neurons, ensuring the similarity measure reflects meaningful changes rather than noise. It also minimizes the impact of irregular spike behavior or augmentation-induced noise in individual channels. Our experiments in subsection 4.2 show that directly comparing entire feature maps leads to ineffective TTA, reinforcing the importance of the per-channel similarity approach. Given that each channel learns distinct features, averaging the similarity within each channel provides a more reliable measure of feature map consistency, improving TTA performance.